Mike has a linear production possibilities frontier that shows maximum production of trucks or computers of 10 million each. Debra also has a linear production possibilities curve that shows maximum production of 9 million trucks or 3 million computers.

Based on the graphs, which of the following statements is true?

Debra has an absolute advantage in the production of computers.
Debra has an absolute advantage in the production of trucks.
Debra has an absolute advantage in the production of computers.
Mike has a comparative advantage in the production of computers.
Answer #1

Maximum production for Mike: 10 million trucks and 10 million computers.

Maximum production for Debra: 9 million trucks and 3 million computers.

Since Mike has greater production in both goods, Mike has an absolute advantage both in the production of trucks as well as that in computers.

However, see that opportunity cost for Mike for production of computers = 10/10 = 1 truck/computer

Opportunity costs for Debra for production of computers = 9/3 = 3 truck/computer.

Hence Mike has more efficiency in making computers and thus he has a comparative advantage in the production of computers.

Correct Ans - D